OpenStudy (anonymous):

convert the following Polar equation into rectangular. r=3sec(theta)

OpenStudy (jdoe0001):

hmmm, I keep on getting a horizontal line :|

OpenStudy (jdoe0001):

but I guess that's it, anyhow, just change around the sec() for its cos() identity :)

OpenStudy (jdoe0001):

a vertical line I meant

OpenStudy (anonymous):

I am still really confused. =(

OpenStudy (jdoe0001):

ok, $$ sec(\theta)=\cfrac{1}{cos(\theta)}\\ so\\ 3sec(\theta) = 3 \times \cfrac{1}{cos(\theta)} \implies \cfrac{3}{cos(\theta)}\\\\ $$

OpenStudy (jdoe0001):

$$ so\\ r=\cfrac{3}{cos(\theta)}\\ \text{now divide both sides by "r", so}\\ 1=\cfrac{3}{rcos(\theta)} \implies rcos(\theta)=\cfrac{3}{1}\\ \text{now keep in mind that } \color{blue}{x = rcos(\theta)} $$

OpenStudy (anonymous):

This is what I got so far, could you help me from here? \[3\times \sqrt{y^2/x^2 +1}=\sqrt{x^2+y^2}\]

OpenStudy (jdoe0001):

see above :) now that's a vertical line :|

OpenStudy (anonymous):

I am so sorry, I am really bad in pre-calculas....I am not getting how to simplify down my answer.

OpenStudy (jdoe0001):

still confused?

OpenStudy (anonymous):

yea I feel really stupid.

OpenStudy (jdoe0001):

well if x = rcos(a); then from the above equations, x=3 which is a vertical line

OpenStudy (anonymous):

so they cancel out leaving only x=3? I get it now. Thank you so much for your help!!

OpenStudy (jdoe0001):